I have 2 problems i need help with,and if you could break it down so i can understand how to solve these.
1:Solve by factoring,
x^2 - 9x = -20
2: Solve by completing the square,
x^2 + 8x + 13 = 0

Thank you

Answer by checkley71(8403)

X~2-9X+20=0 OR (X-4)(X-5)=0 OR X=4 & X=5
X~2+8X=-13 OR X~2+8X+16=-13+16 OR (X+4)(X+4)=3 OR X+4=SQRT3 OR X=SQRT3-4